# Sequential Monte Carlo Squared for Online Inference in Stochastic Epidemic Models

### v1

@@ 2.3.2 SMC2 scheme
Given a weighted sample $\{\theta^{m}_{t},\omega^{m}_{t}\}_{m=1}^{N_{\theta}}$ from $p(\theta|y_{1:t})$, the SMC2 algorithm reweights each $\theta$-particles according to an unbiased estimate of the incremental likelihood $p(y_{t}|y_{1:t-1},\theta^{m}_{t})$, obtained from the particle filter. As time progresses, there is a risk that only a few $\theta$-particles will accumulate significant weight. To address this, a resampling step is performed to generate a new set of particles from the dominant ones, followed by a PMCMC move step, such as the particle Marginal Metropolis-Hastings (PMMH) kernel, which preserves the target posterior distribution (Andrieu et al., 2010). The combination of the resampling and move steps is referred to as the “rejuvenation” step. In this study, new parameters are proposed from a normal distribution $\theta^{*}\sim q(\theta^{*}|\theta^{m}_{t}):=\mathcal{N}(\hat{\mu}_{t},c\hat{\Sigma}_{t})$, where $c$ is a scaling factor. The mean and variance-covariance matrix are given by:
$\hat{\mu}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\theta^{m}_{t},\text{ and }\hat{\Sigma}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}(\theta^{m}_{t}-\hat{\mu}_{t})(\theta^{m}_{t}-\hat{\mu}_{t})^{T},$ (10)
@@
4:     Update importance weights: $\omega^{m}_{t}=\omega^{m}_{t-1}\hat{p}_{N_{x}}(y_{t}|y_{1:t-1},\theta^{m}_{t}),\quad\omega^{m}_{t}=\omega^{m}_{t}/\sum_{j=1}^{N_{x}}\omega^{j}_{t}$
5:     if ESS $<\tau_{R}N_{\theta}$ then
@@
where $\hat{p}_{N_{x}}(y_{s}|y_{1:s-1},\theta^{m}_{s})$ is the likelihood computed using the particle filter, as described in equation (8). Each rejuvenation (step 8-12 of the Algorithm 2) performs $M$ successive PMMH moves on each particle $\theta^{m}_{t}$. The resampling step is similar to the one used in Algorithm 1 but is performed only when a degeneracy criterion is met. A common way is to estimate the degree of degeneracy by computing the effective sample size (ESS) :
$\text{ESS}=\left(\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\right)^{2}\bigg{/}\sum_{m=1}^{N_{\theta}}(\omega^{m}_{t})^{2}.$ (12)
The rejuvenation step is triggered if the ESS falls below a specified threshold, typically if ESS $<\tau_{R}N_{\theta}$, where $\tau_{R}\in(0,1)$ is a user-defined parameter. The choice of $\tau_{R}$ controls the frequency of rejuvenation steps. A lower threshold for ESS leads to more frequent rejuvenation, which increases the computational cost. In practice, an ESS threshold of 50% ($\tau_{R}=0.5$) is commonly used (Chopin et al., 2013; Jacob, 2015).
@@ 2.4 Online SMC2
While SMC2 allows for sequential estimation of parameters as new observations arrive, it is not a truly online method. The PMMH step, which has a computational complexity of $\mathcal{O}(tN_{\theta}N_{x})$, involves a calculation of the total likelihood. This complexity increases with the number of observations $t$, making it computationally expensive as more data becomes available. Although Chopin et al. (2013) demonstrates that, in theory, the frequency at which the rejuvenation step is performed will decrease over time, the algorithm may still be impractical for some real-time applications. A possible solution suggested by Vieira (2018) to implement SMC2 in an online manner (O-SMC2) is to restrict the likelihood calculation set to a fixed-size subset of past observations, keeping the computational cost bounded. By considering a fixed window of observations of size $t_{k}>0$ in the PMMH, we can ensure that the online part of SMC2 will not be altered, and the online approximation will target the posterior of the parameters given the observation window $p(\theta|y_{t-t_{k}:t})$. This approach changes the cost from $\mathcal{O}(tN_{\theta}N_{x})$ to $\mathcal{O}(t_{k}N_{\theta}N_{x})$.
@@
$\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{m}_{t})p(\theta^{m}_{t})}\times\dfrac{q(\theta^{m}_{t}|\theta^{*})}{q(\theta^{*}|\theta^{m}_{t})}\right\},$ (13)
where the estimate of the marginal likelihood within the time window is computed similarly to (9), but this time using the restricted particle set $x_{t-t_{k}:t}^{1:N_{x}}$ ( see A for further details) :
$\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right).$ (14)
@@ 3.1.1 Experiment 1
Figure 3 displays the sequential inference of daily new cases of infection together with the static parameters. We can observe that the predicted new cases are in good agreement with the data. The results also indicate that all true parameter values fall within the 50% credible interval, closely aligning with the posterior median as we evolve over time. Initially, during the epidemic’s onset, the data lacks sufficient information within the time interval $[0,20]$, resulting in widespread parameter uncertainty. However, as we get closer to the peak of the epidemic, the inference algorithm captures the dynamics of the epidemic, notably reducing parameter uncertainty. As the dynamics begin to slow around time $t=60$, the posterior uncertainty stays about the same.
@@ 3.1.2 Experiment 2
In this experiment, we introduce variability in the transmission rate to simulate a more realistic scenario. Synthetic data are generated using a time-varying $\beta(t)$, modeled as:
@@
To accommodate the non-constant transmission rate, we model $\beta_{t}$ in the model (2.1) as a geometric Brownian motion:
$\beta_{t}=e^{\log(\beta_{t-1})+\varepsilon_{t}},\quad\varepsilon_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (16)
[…]
Figure 4: State and parameter inference for Experiment 2. Top row: Left - Filtering estimates of daily new infections with observed values in orange; Right - Filtering estimate of the time-dependent transmission rate $\beta_{t}$. Middle row: Filtering estimates of $\sigma$, $\gamma$, and $\nu_{\beta}$ over time. Bottom row: Posterior distributions of $\sigma$, $\gamma$, and $\nu_{\beta}$ the final time step, with 95% credible intervals in dashed blue lines. In rows 1 and 2, the solid black lines show the median estimate; shaded blue regions indicate posterior uncertainty quantiles. True parameter values are represented by orange dashed lines.
Figure 4 demonstrates the ability of the method to track multiple outbreak waves and capture the trend in a time-varying transmission rate, albeit with a minor lag. This slight delay is expected, as the analysis relies on daily case data (i.e., the number of newly infected individuals transitioning from $E$ to $I$) rather than exact counts of new infections (i.e., individuals moving from $S$ to $E$), which are often unknown during an epidemic. As with Experiment 1, the true parameter values lie within the 50% credible interval, indicating reliable parameter estimation even under conditions of variable transmission.
To evaluate the performance of the O-SMC2 algorithm, we tested a range of window sizes $t_{k}$, keeping the priors and number of particle configurations consistent across experiments. Tables B.1 and B.2 in C show that even with smaller window sizes, the algorithm produces reasonable parameter estimates that are close to the true values, with credible intervals comparable to those of standard SMC2 (although slightly larger for smaller window sizes). These findings align with previous research (Vieira (2018), Chapter 15). Figures B.1-B.2 display the posterior densities of the parameters at the final time step, demonstrating consistency with estimates obtained using SMC2.
@@ 3.2 Real case study: Analysis of COVID-19 in Ireland
Since the onset of the COVID-19 pandemic in December 2019 in Wuhan, China, the World Health Organization (WHO) has reported over 772 million confirmed cases globally and more than 69 million deaths (World Health Organisation (2024), WHO). The rapid transmission and widespread nature of the virus led the WHO to declare it a pandemic. In (Cazelles et al., 2021), particle MCMC has been employed to reconstruct the COVID-19 epidemic’s trajectory in Ireland. However, it is worth noting that while this method efficiently tracks observed data, it remains an offline approach and may not be suitable for real-time decision-making when rapid responses are required.
@@ 3.2.1 Model
We present a fully discrete stochastic extended-SEIR model to study the spread of COVID-19, with a structure tailored to focus on the change in transmission dynamics during the pandemic. In this model, individuals transition through compartments representing different stages of infection. The model comprises the following compartments: $S_{t}$ for susceptible individuals, $E_{t}$ for exposed but not yet infectious individuals, $A_{t}$ for asymptomatic infected individuals, $I_{t}$ for symptomatic infected individuals, and finally, $R_{t}$ is the count of individuals recovering or succumbing to the virus. The following discrete equations describe the dynamics of the model:
@@
where $Y_{AB}(t)$ represents the number of individuals moving from compartment $A$ to compartment $B$ at time $t$ and defined by the following binomial distributions:
$Y_{SE}(t)\sim\mathrm{Bin}\left(S_{t},1-\exp\left(-\beta_{t}\frac{(I_{t}+r_{A}A_{t})}{N}\cdot\delta t\right)\right),$
$Y_{E}(t)\sim\mathrm{Bin}\left(E_{t},1-e^{-\sigma\cdot\delta t}\right),\quad Y_{EA}(t)\sim\mathrm{Bin}\left(Y_{E}(t),p_{A}\right),$ (18)
$Y_{AR}(t)\sim\mathrm{Bin}\left(A_{t},1-e^{-\gamma\cdot\delta t}\right),\quad Y_{IR}(t)\sim\mathrm{Bin}\left(I_{t},1-e^{-\gamma\cdot\delta t}\right).$
@@
The models presented in (3.2.1)-(3.2.1), do not explicitly incorporate vaccination. This is because they are primarily designed to focus on the spread of the virus and the impact of various non-pharmaceutical interventions during a period when vaccines are either unavailable or not yet widely deployed. However, the effect of vaccination can easily be introduced by adding a compartment for vaccinated individuals and adjusting the transmission dynamics to reflect partial or complete immunity. Allowing the transmission rate to be time-varying seems reasonable since, throughout the outbreak, the intensity of disease transmission will tend to vary due to individuals’ behavioral changes or government lockdown measures. Hence, we consider a geometric random walk (RW) model for the transmission rate:
$\beta_{t}=e^{\log(\beta_{t-1})+w_{t}},\quad w_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (20)
where $\nu_{\beta}$ is the parameter controlling the innovation in the transmission rate, consistent with previous study of COVID-19 in Ireland (Cazelles et al., 2021). Giving the distribution of $\beta_{t}$, we can sequentially estimate the distribution of the effective reproduction number ($R_{\text{eff}}(t)$) that refers to the average number of secondary infections caused by a single infected individual during their infectious period in a population where some individuals are immune or other interventions have been implemented:
$R_{\text{eff}}(t)=\beta_{t}\left(\frac{(1-p_{A})+p_{A}\cdot r_{A}}{\gamma}\right).\frac{S_{t}}{N}.$ (21)

### v2

@@ 2.3.2 SMC2 scheme
Given a weighted sample $\{\theta^{m}_{t},\omega^{m}_{t}\}_{m=1}^{N_{\theta}}$ from $p(\theta|y_{1:t})$, the SMC2 algorithm reweights each $\theta$-particle using an unbiased estimate of the incremental likelihood $p(y_{t}|y_{1:t-1},\theta^{m}_{t})$, obtained from the Bootstrap particle filter. Alternative estimates from other filters, such as the auxiliary particle filter, can also be employed to improve efficiency (Golightly and Kypraios, 2018). As we accrue observations through time, there is a risk that only a small number of $\theta$-particles will hold a significant proportion of the weight. To address this, a resampling step is performed to generate a new set of particles from the dominant ones, followed by a PMCMC move step, such as the particle Marginal Metropolis-Hastings (PMMH) kernel, which preserves the target posterior distribution (Andrieu et al., 2010). The combination of the resampling and move steps is referred to as the “rejuvenation” step. In this study, new parameters are proposed from a normal distribution $\theta^{*}\sim q(\theta^{*}|\theta^{m}_{t}):=\mathcal{N}(\hat{\mu}_{t},c\hat{\Sigma}_{t})$, where $c$ is a scaling factor. The mean and variance-covariance matrix are given by:
$\hat{\mu}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\theta^{m}_{t},\text{ and }\hat{\Sigma}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}(\theta^{m}_{t}-\hat{\mu}_{t})(\theta^{m}_{t}-\hat{\mu}_{t})^{\top},$ (10)
@@
where $\hat{p}_{N_{x}}(y_{s}|y_{1:s-1},\theta^{m}_{s})$ is the likelihood computed using the particle filter, as described in equation (8). Each rejuvenation (step 8-12 of the Algorithm 2) performs $M$ successive PMMH moves on each particle $\theta^{m}_{t}$. The resampling step is similar to the one used in Algorithm 1 but is performed only when a degeneracy criterion is met.
A common way is to estimate the degree of degeneracy by computing the effective sample size (ESS) :
$\text{ESS}_{\theta}=\left(\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\right)^{2}\bigg{/}\sum_{m=1}^{N_{\theta}}(\omega^{m}_{t})^{2}.$ (12)
[…]
@@
4:  Update importance weights: $\omega^{m}_{t}=\omega^{m}_{t-1}\hat{p}_{N_{x}}(y_{t}|y_{1:t-1},\theta^{m}_{t}),\quad\omega^{m}_{t}=\omega^{m}_{t}/\sum_{j=1}^{N_{\theta}}\omega^{j}_{t}$
[…]
@@ 2.4 Online SMC2
While SMC2 enables sequential estimation of parameters as new observations arrive, it is not a truly online method. The PMMH step, which has a computational complexity of $\mathcal{O}(tN_{\theta}N_{x})$, requires calculating the marginal likelihood $p(y_{1:t}|\theta)$. Consequently, the computational cost increases with the number of observations $t$, making the method increasingly expensive as more data accumulates. Although Chopin et al. (2013) demonstrates that, in theory, the frequency at which the rejuvenation step is performed will decrease over time, the algorithm may still be impractical for some real-time applications. Vieira (2018) proposes to implement SMC2 in an online manner (O-SMC2) by calculating the likelihood using only a fixed-size window of past observations. This restriction keeps the computational cost of a likelihood calculation constant. By considering a fixed window of observations of size $t_{k}>0$ in the PMMH, we can ensure that the online part of SMC2 will not be altered, and the online approximation will target the posterior of the parameters given the observation window $p(\theta|y_{t-t_{k}+1:t})$. This reduces the computational cost of each PMMH iteration from $\mathcal{O}(tN_{\theta}N_{x})$ to $\mathcal{O}(t_{k}N_{\theta}N_{x})$.
@@
$\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta^{m}_{t})p(\theta^{m}_{t})}\times\dfrac{q(\theta^{m}_{t}|\theta^{*})}{q(\theta^{*}|\theta^{m}_{t})}\right\},$ (13)
where the estimate of the marginal likelihood within the time window is computed similarly to (9), but this time using the restricted particle set $x_{t-t_{k}+1:t}^{1:N_{x}}$ (see B for further details) :
$\hat{p}_{N_{x}}(y_{t-t_{k}+1:t}|\theta)=\prod_{s=t-t_{k}+1}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right).$ (14)
@@
Figure 3: Flowchart of the SMC2/O-SMC2 algorithm.
Figure 3 illustrates the flow of the SMC2/O-SMC2 algorithm. At each time step, a new observation $y_{t}$ arrives. A particle filter is run for each parameter particle $\theta^{m}$ to estimate the likelihood and update the weights. If resampling is needed, it is followed by a PMMH rejuvenation step. The procedure repeats as new observations arrive, allowing for sequential joint inference of parameters and latent states.
2.5 Structural identifiability
Prior to presenting the inference results using the O-SMC2 methodology in Section 3, we conduct a structural identifiability analysis of the SEIR model parameters. This analysis employed a deterministic mean-field approximation (Roberts et al., 2015) of the stochastic model given in Equation (2.1) and was performed using the StructuralIdentifiability.jl package in Julia (Dong et al., 2023; Liyanage et al., 2025). The results indicate that the parameters $\beta$, $\sigma$, and $\gamma$ are globally identifiable from ideal observations of the daily incidence. This provides a theoretical basis supporting the feasibility of parameter estimation under perfect data conditions. Further details of the model formulation and the output function used are provided in Appendix A.
However, it is important to emphasize some limitations inherent in structural identifiability analysis. The analysis makes the assumptions of noise-free, continuously reported and complete observations. These assumptions will rarely hold in practical epidemiological settings. Epidemiological data are often sparse, noisy, and collected at discrete time points (for example, daily or weekly reports), and may have missing information. These challenges are especially pronounced in the early stages of an outbreak.
In the context of sequential parameter estimation, when only a small number of observations have been accrued at the early stages of an outbreak, we expect to see high uncertainty, wide credible intervals, and possible parameter confounding (Sauer et al., 2021). Thankfully, this is usually observed only during early phases. With the incorporation of informative priors, these challenges are somewhat mitigated and due to the adaptive qualities of sequential inference, they gradually resolve as we accrue more data. We may also experience challenges during periods of abrupt change in disease dynamics. Parameters may become highly correlated and hence poorly identified. This again can be mitigated through the use of informative priors.
In summary, the conclusions drawn from an identifiability analysis are subject to often unrealisable assumptions in practical settings. In our context, idealized structural analyses should be viewed as necessary but not sufficient conditions for practical identifiability.
@@ 3.1.1 Experiment 1
Figure 4 displays the sequential inference of daily new cases of infection together with the static parameters. We can observe that the predicted new cases are in good agreement with the data. The results also indicate that all true parameter values fall within the 95% credible interval, closely aligning with the posterior median as we evolve over time. Initially, during the epidemic’s onset, the data lacks sufficient information within the time interval $[0,20]$, resulting in widespread parameter uncertainty. However, as the epidemic approaches its peak, the inference algorithm captures the dynamics well, notably reducing parameter uncertainty. As the dynamics begin to slow around time $t=60$, the posterior uncertainty stabilizes.
@@ 3.1.2 Experiment 2
In this experiment, we introduce variability in the transmission rate to simulate a more realistic scenario. Synthetic data are generated using a time-varying $\beta=\beta(t)$, modeled as:
@@
To accommodate the non-constant transmission rate, we model $\beta(t)$ in the model (2.1) as a geometric Brownian motion:
$\beta(t)=e^{\log(\beta(t-1))+\varepsilon_{t}},\quad\varepsilon_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (16)
where $\nu_{\beta}$ governs the variability in the transmission rate. The latent state vector is now $x_{t}=(S_{t},E_{t},I_{t},R_{t},\beta(t))$, and the parameters to estimate are $\theta=(\sigma,\gamma,\nu_{\beta})$. The O-SMC2 algorithm uses the same initial particle state as in Experiment 1, with $\beta(0)\sim\mathcal{U}(0.2,0.5)$.
Figure 5: State and parameter inference for Experiment 2. Top row: Left - Filtering estimates of daily new infections with observed values in orange; Right - Filtering estimate of the time-dependent transmission rate $\beta(t)$. Middle row: Filtering estimates of $\sigma$, $\gamma$, and $\nu_{\beta}$ over time. Bottom row: Posterior distributions of $\sigma$, $\gamma$, and $\nu_{\beta}$ the final time step, with 95% credible intervals in dashed blue lines. In rows 1 and 2, the solid black lines show the median estimate; shaded blue regions indicate posterior uncertainty quantiles. True parameter values are represented by orange dashed lines.
Priors for the parameters are $\sigma\sim\mathcal{U}(0.3,0.6)$, $\gamma\sim\mathcal{TN}_{[0,1]}(0.14,0.01)$, and $\nu_{\beta}\sim\mathcal{TN}_{[0.05,0.15]}(0.1,0.05)$. A more restrictive prior on $\gamma$ is used due to its high correlation with the transmission rate $\beta(t)$, especially during abrupt changes in transmission dynamics. Fluctuations in new cases can reflect changes in $\beta(t)$, but could also be influenced by $\gamma$ as shown in (Inouzhe et al., 2023). The prior for $\nu_{\beta}$ allows for rapid changes in the transmission rate, ensuring the model can adapt to sudden shifts in disease dynamics and aligns with values observed in the literature (Funk et al., 2018; Calvetti et al., 2021).
Figure 5 demonstrates the ability of the method to track multiple outbreak waves and capture the trend in a time-varying transmission rate, albeit with a minor lag. This slight delay is expected, as the analysis relies on daily case data (i.e., the number of newly infectious individuals transitioning from $E$ to $I$) rather than exact counts of new infections (i.e., individuals moving from $S$ to $E$), which are often unknown during an epidemic. As with Experiment 1, the true parameter values lie within the 95% credible interval, indicating reliable parameter estimation even under conditions of variable transmission.
To evaluate the performance of the O-SMC2 algorithm, we tested a range of window sizes $t_{k}$, keeping the priors and number of particle configurations consistent across experiments. Tables 2 and 3 in Appendix C show that even with smaller window sizes, the algorithm produces reasonable parameter estimates that are close to the true values, with credible intervals comparable to those of standard SMC2 approach (i.e., when all past observations are used at each time step), although slightly larger for smaller window sizes. Notably, the computational cost decreases substantially with smaller windows, making O-SMC2 well-suited for real-time or resource-limited settings. These findings align with previous research (Vieira (2018), Chapter 15). Figures 1-2 in Appendix C display the posterior densities of the parameters at the final time step, demonstrating consistency with estimates obtained using SMC2.
@@ 3.2 Real case study: Analysis of COVID-19 in Ireland
Since the onset of the COVID-19 pandemic in December 2019 in Wuhan, China, the World Health Organization (WHO) has reported over 772 million confirmed cases globally and more than 69 million deaths (World Health Organisation (2024), WHO). The rapid transmission and widespread nature of the virus led the WHO to declare it a pandemic. In Cazelles et al. (2021), particle MCMC has been employed to reconstruct the COVID-19 epidemic’s trajectory in Ireland. However, it is worth noting that while this method efficiently tracks observed data, it remains an offline approach and may not be suitable for real-time decision-making when rapid responses are required.
@@ 3.2.1 Model
We present a fully discrete stochastic extended-SEIR model to study the spread of COVID-19, with a structure designed to capture changes in transmission dynamics during the pandemic. In this model, individuals transition through compartments representing different stages of infection: $S_{t}$ representing susceptible individuals, $E_{t}$ for those who are exposed and infected but not yet infectious, $A_{t}$ for asymptomatic infectious individuals, $I_{t}$ for symptomatic infectious individuals, and finally, $R_{t}$ is the number of individuals recovering or succumbing to the virus. The following discrete equations describe the dynamics of the model:
@@
where $Y_{AB}(t)$ represents the number of individuals moving from compartment $A$ to compartment $B$ at time $t$ and these quantities are defined through the binomial random variables:
$Y_{SE}(t)\sim\mathrm{Bin}\left(S_{t},1-\exp\left(-\beta(t)\frac{(I_{t}+r_{A}A_{t})}{N}\delta t\right)\right),$
$Y_{E}(t)\sim\mathrm{Bin}\left(E_{t},1-e^{-\sigma\delta t}\right),\quad Y_{EA}(t)\sim\mathrm{Bin}\left(Y_{E}(t),p_{A}\right),$ (18)
$Y_{AR}(t)\sim\mathrm{Bin}\left(A_{t},1-e^{-\gamma\delta t}\right),\quad Y_{IR}(t)\sim\mathrm{Bin}\left(I_{t},1-e^{-\gamma\delta t}\right).$
@@
The models presented in (3.2.1)-(3.2.1), do not explicitly incorporate vaccination. This is because they are primarily designed to focus on the spread of the virus and the impact of various non-pharmaceutical interventions during a period when vaccines are either unavailable or not yet widely deployed. However, the effect of vaccination can easily be introduced by adding a compartment for vaccinated individuals and adjusting the transmission dynamics to reflect partial or complete immunity. Allowing the transmission rate to be time-varying seems reasonable since, throughout the epidemic, the intensity of disease transmission will tend to vary due to individuals’ behavioral changes or government lockdown measures. Hence, we consider a geometric random walk (RW) model for the transmission rate:
$\beta(t)=e^{\log(\beta(t-1))+w_{t}},\quad w_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (20)
where $\nu_{\beta}$ is the parameter controlling the innovation in the transmission rate, consistent with previous study of COVID-19 in Ireland (Cazelles et al., 2021). Giving the distribution of $\beta(t)$, we can sequentially estimate the distribution of the effective reproduction number ($R_{\text{eff}}(t)$) that refers to the average number of secondary infections caused by a single infected individual during their infectious period in a population where some individuals are immune or other interventions have been implemented:
$R_{\text{eff}}(t)=\beta(t)\left(\frac{(1-p_{A})+p_{A}r_{A}}{\gamma}\right)\frac{S_{t}}{N}.$ (21)